# A topological counting rule for shells

### v1

@@ Abstract
Holding a shell in their hands, one can apply six loads: three by pulling and shearing, and three by bending and twisting. Here, it is shown that the shell will resist exactly three load cases and comply with the other three, provided the shell is simply connected, meaning it has no holes and no handles.
@@
Here, a new counting rule is established for simply connected shells and trusses that discretize them. These trusses are locally isostatic but topologically trivial in the Kane–Lubensky sense. Nonetheless, the simply connected topology enables a count of macroscopic isometric deformations, i.e., membrane-like or flexure-like deformations that are uniform up to a certain correction. The count is simple: there exist exactly three such isometric deformations. Conversely, a shell resists exactly three macroscopic loads, tension and bending combined, provided it is simply connected, irrespective of its composition or geometry. Note that for modes to be meaningfully classified as membrane, flexure, or macroscopic, the truss must be periodic or statistically homogeneous. If not, a version of the rule holds still but is not as amenable to interpretation. The proof relies on two familiar results, seldom used in tandem: the static-geometric analogy, and the Hill-Mandel lemma. Both are introduced next, followed by the proof and by a discussion of the role of topology.
Figure 1: The static-geometric analogy for macroscopic fields. Wide arrows are admissible stresses; thin arrows are isometric deformations: (a) in a plane, admissible longitudinal tension corresponds to isometric lateral bending; (b) in a singly corrugated membrane, an admissible longitudinal moment corresponds to isometric lateral extension; (c) in a multiply connected grid of tubes, admissible longitudinal tension has no corresponding isometric lateral bending.
Consider a shell with a midsurface $z=z(x_{1},x_{2})$, of Hessian $\mathbf{H}_{z}$, and with a statically admissible distribution of membrane stresses $\bm{\sigma}$, namely such that
$\operatorname{div}\bm{\sigma}=0,\quad\mathbf{H}_{z}:\bm{\sigma}=0.$ (1)
Letting $\bm{\sigma}=\operatorname{cof}\mathbf{H}_{\phi}$ derive from an Airy stress function $\phi$ eliminates the first equation and recasts the second into Pucher’s equation [24]
@@
Now if a displacement $(\mathbf{u}\equiv(u_{1},u_{2}),w)$ is isometric, then the infinitesimal strain
@@
vanishes meaning that $\bm{\nabla}w\otimes^{s}\bm{\nabla}z$ is a compatible planar strain. The compatibility equation, expanded and reduced, turns out to be identical to (2) with $w$ replacing $\phi$. This is the essence of the static-geometric analogy: a deflection is isometric if, as a stress function, it is statically admissible, and conversely. The analogy is well-known [25]; here, the novelty shall be in completing it with integral formulas for average strains and stresses. In particular, for a statistically homogeneous, say periodic, shell let
@@
be the displacement field due to a macroscopic membrane strain $\mathbf{E}$ and a macroscopic bending strain $\bm{\chi}$. If the deformation is isometric, then
@@
where $\left\langle{}\right\rangle$ averages over the 2D unit cell, or a representative surface element in the stochastic case. The subscript ‘$o$’ will be used to denote isometric strains and will be dropped when $\mathbf{E}$ and $\bm{\chi}$ are meant to be general strains. Similarly, for a statically admissible periodic membrane stress $\bm{\sigma}$ of Airy stress $\phi$, the macroscopic membrane and bending stresses are
@@
Thus, the static-geometric analogy further matches $\operatorname{cof}\bm{\chi}$ and $\bm{\Sigma}$ as well as $\operatorname{cof}\mathbf{E}_{o}$ and $\mathbf{M}_{o}$: overall, it provides an isomorphism
@@
between the spaces of macroscopic isometric deformations and of macroscopic admissible stresses, where the $\operatorname{cof}$ operator plays the role of a $90^{\circ}$ rotation; see Figure 1.
The other main ingredient is the Hill-Mandel lemma [26]. This is a useful result of homogenization theory that describes some boundary conditions such that the virtual stress-strain work is the same if computed using the local fields or the macroscopic fields. Namely, for any admissible periodic $\bm{\sigma}$ and any admissible $(\mathbf{u},w)$ in the sense of (4), one has
@@
The proof is straightforward and employs the divergence theorem in conjunction with periodicity (4), admissibility (1) and the compatibility conditions (3).
What is lacking here, to apply arguments à la Maxwell-Calladine, is to identify the spaces of macroscopic admissible stresses and isometric strains with the image and kernel of some structural matrix. In the present context, the appropriate matrix turns out to be the matrix of effective elasticity moduli $\mathbf{A}$. Define it by the minimization procedure
$\begin{bmatrix}\mathbf{E}\\ \bm{\chi}\end{bmatrix}^{T}\mathbf{A}\begin{bmatrix}\mathbf{E}\\ \bm{\chi}\end{bmatrix}\equiv\min_{(\mathbf{u},w)}\left\langle{\bm{\varepsilon}:\mathbf{c}:\bm{\varepsilon}}\right\rangle$ (9)
under the constraints of equation (4), where $\mathbf{c}$ is the shell’s membrane stiffness tensor. The goal is to show that the spaces $\text{Im}\mathbf{A}$ and $\text{Ker}\mathbf{A}$ are precisely the spaces of admissible macroscopic stresses and isometric macroscopic strains.
It is rather clear that a macroscopic strain in the kernel of $\mathbf{A}$ is isometric for it has zero energy and therefore zero membrane strain. For the converse, the polarization identity of the bilinear form $\mathbf{A}$ is used: a macroscopic isometric strain $(\mathbf{E}_{o},\bm{\chi}_{o})$ can be superposed to an arbitrary strain $(\mathbf{E},\bm{\chi})$ without changing the strain energy, since the local membrane strain does not change either. By taking the difference, it comes that
$\begin{bmatrix}\mathbf{E}\\ \bm{\chi}\end{bmatrix}^{T}\mathbf{A}\begin{bmatrix}\mathbf{E}_{o}\\ \bm{\chi}_{o}\end{bmatrix}=0,\text{ for all }(\mathbf{E},\bm{\chi}),$ (10)
meaning that macroscopic isometric strains are in the kernel of $\mathbf{A}$. As for stresses, let $(\bm{\Sigma},\mathbf{M})$ be in the image of $\mathbf{A}$ produced by some $(\mathbf{E},\bm{\chi})$. Let $\bm{\sigma}=\mathbf{c}:\bm{\varepsilon}$ be the membrane stresses obtained by the minimization of strain energy (9). The stationarity conditions readily imply that $\bm{\sigma}$ is statically admissible and the Hill-Mandel lemma ensures that $\bm{\Sigma}=\left\langle{\bm{\sigma}}\right\rangle$ and $\mathbf{M}=\left\langle{z\bm{\sigma}}\right\rangle$. Conversely, let $(\bm{\Sigma},\mathbf{M})$ be an admissible macroscopic stress obtained by averaging an admissible periodic membrane stress $\bm{\sigma}$. Testing $\bm{\sigma}$ against isometric strains shows, by the Hill-Mandel lemma, that $(\bm{\Sigma},\mathbf{M})$ is in the orthogonal to the kernel of $\mathbf{A}$ and therefore, by symmetry of $\mathbf{A}$, in the image of $\mathbf{A}$.
In conclusion, $\text{Im}\mathbf{A}\oplus\text{Ker}\mathbf{A}$ provides a decomposition of the space of pairs of symmetric second order tensors into a direct sum of the spaces of admissible macroscopic stresses and isometric macroscopic strains, while the analogy previously showed that these spaces are isomorphic:
$\begin{gathered}\{(\mathbf{\Sigma},\mathbf{M})\}\oplus\{(\mathbf{E}_{o},\bm{\chi}_{o})\}=S^{2}\times S^{2},\\ \{(\mathbf{\Sigma},\mathbf{M})\}\cong\{(\mathbf{E}_{o},\bm{\chi}_{o})\},\end{gathered}$ (11)
where $S^{2}$ is the 3D space of symmetric second order tensors. Taking the dimensions yields at last the counting rule
$\dim\{(\mathbf{E}_{o},\bm{\chi}_{o})\}=3.$ (12)
The topology of the shell plays a crucial role that remained implicit in the proof presented above and that should be highlighted. On one hand, suppose the shell had handles, then there would be topological obstructions to integrating an admissible stress $\bm{\sigma}$ into an Airy stress function $\phi$. The static-geometric analogy, for averages, will then only encapsulate states of stress that derive from $\phi$ meaning that the rule needs to be weakened into
$\dim\{(\mathbf{E}_{o},\bm{\chi}_{o})\}\leq 3\quad\text{(with handles)}.$ (13)
On the other hand, suppose the shell had holes, then there would be extra admissibility constraints for stresses in the form of zero tractions at boundaries. The analogy fails by leaving out isometries that do not reproduce this condition, that is isometries that cannot be extended over the holes or, equivalently, that deform the holes. The rule then weakens into
$\dim\{(\mathbf{E}_{o},\bm{\chi}_{o})\}\geq 3\quad\text{(with holes)}.$ (14)
Through the involved isomorphism, the rule endows the space of macroscopic isometric deformations with a natural symplectic structure: for any pair of modes $(\mathbf{E}_{i},\bm{\chi}_{i})_{i=1,2}$ in that space
@@
This is a restatement of the Hill-Mandel lemma (8) with the isomorphism (7) in mind. Such a relationship was recently derived using purely geometric tools [27, 28]. Here, the novelty is in the topological counting rule. The above identity reduces, when either mode is a pure membrane, e.g., $\bm{\chi}_{1}=\mathbf{0}$, or a pure flexure mode, e.g., $\mathbf{E}_{1}=\mathbf{0}$, to a remarkable identity between the in- and out-of-plane effective Poisson’s coefficient of the shell [29, 30, 31, 32, 33, 34]. Thanks to the construction of the analogy-induced isomorphism, it is possible to draw a whole series of exact relationships among the effective moduli, all summarized into
@@
This relationship is more robust than the counting rule strictly speaking. Indeed, one can reasonably assume that the effective tensor $\mathbf{A}$ depends continuously on the local membrane stiffness $\mathbf{c}$ which can then be taken to 0 at places so as to create holes. By contrast, the above relationship can certainly fail in the presence of handles, in which case $\mathbf{A}$ can become full rank (e.g., as in a foldcore sandwich panels [35]).
@@
Examples are abundant where a shell is flexible out of plane and stiff in the plane, including trivially the plane, but also pyramidal trusses [40] and the Ron Resch origami pattern [41]. There are also many examples of shells with a unique pure membrane mode (i.e., with $\bm{\chi}_{o}=\mathbf{0}$) and two flexure modes including the Miura ori, the eggbox pattern and any singly corrugated shell [20]. Beyond that, there are no reported examples of shells with two or three pure membrane modes. It is not known if there is a fundamental limit on the number of pure membrane modes, but this appears to be unlikely. Here, the unit cell of a shell with approximately three pure membrane modes is produced by numerical minimization of the trace of the effective membrane stiffness. While not a proof, the numerical experiment convincingly shows that it is possible for a periodic simply connected membrane to be stiff out of plane. The simulations employ a truss code with periodic boundary conditions to compute the effective membrane stiffness along with a standard minimization routine to minimize its trace. See  [^1] Code for membrane stiffness minimization is available at https://github.com/nassarh/ for implementation details; see also [43]. Figure 2 shows the results. Therein, it is remarkable that the optimized geometries achieve a reduction of the effective membrane stiffness by 10 orders of magnitude, with as little as $20\%$ changes in elevations, and no noticeable qualitative changes in the geometry. This reinforces the point of view that, however insightful, the present methods leave important quantitative questions unanswered, e.g., regarding the sensitivity of the space of isometries to perturbations.
Figure 2: Minimization of the effective membrane stiffness for discrete shells with (a) $2\times 2$, (b) $4\times 4$, and (c) $6\times 6$ nodes per unit cell. Random initial geometries are in the top row; geometries with optimized elevations are in the bottom row. The trace of the effective membrane stiffness tensor is reduced by a factor of $\sim 2$, of $\sim 10^{10}$ and of $\sim 10^{10}$, respectively.

### v2

@@ Abstract
Holding a shell in their hands, one can apply six loads: three by pulling and shearing, and three by bending and twisting. Here, it is shown that the shell resists exactly three load cases and comply with the other three, provided the shell is simply connected, meaning it has no holes and no handles. Formally, it is shown that the space of homogeneous membrane and bending strains, defined in the sense of plate theory, that can be relaxed into an infinitesimal isometry by a periodic, or a statistically homogeneous, deflection is three-dimensional for any simply-connected periodic, or statistically homogeneous, shell, be it corrugated, creased or wrinkled.
@@
Figure 1: The static-geometric analogy for effective fields. Wide arrows are admissible stresses; thin arrows are isometric deformations: (a) in a plane, admissible longitudinal tension corresponds to isometric lateral bending; (b) in a singly corrugated membrane, an admissible longitudinal moment corresponds to isometric lateral extension; (c) in a multiply connected grid of tubes, admissible longitudinal tension has no corresponding isometric lateral bending.
Here, a new counting rule is established for simply connected shells and trusses that discretize them. These trusses are locally isostatic but topologically trivial in the Kane–Lubensky sense. Nonetheless, the simply connected topology enables a count of effective isometric deformations, i.e., membrane-like or flexure-like deformations that are uniform up to a certain correction. The count is simple: there exist exactly three such isometric deformations. Conversely, a shell resists exactly three effective loads, tension and bending combined, provided it is simply connected, irrespective of its composition or geometry. Note that for modes to be meaningfully classified as membrane, flexure, or effective, the truss must be periodic or statistically homogeneous. The proof relies on two familiar results, seldom used in tandem: the static-geometric analogy, and the Hill-Mandel lemma. Both are introduced under the assumptions of smooth periodic graphs along with a proof of the counting rule. A discussion of the role of topology follows and the validity of the proof for more general, piecewise-smooth and/or non-periodic, shells is discussed.
The static-geometric analogy
Consider a shell whose midsurface is the graph of a function $z=z(x_{1},x_{2})$. A statically admissible distribution of membrane stresses $\bm{\sigma}$ satisfies two force balances, namely
$\operatorname{div}\bm{\sigma}=0,\quad\mathbf{H}_{z}:\bm{\sigma}=0,$ (1)
where $\mathbf{H}_{z}$ is the Hessian of $z$ and the colon denotes a double contraction. The first is force balance of surface tensions (i.e., in the plane tangent to the shell); the second is force balance in the normal direction, identical to the well-known Laplace law (minus the internal pressure). Letting $\bm{\sigma}=\operatorname{cof}\mathbf{H}_{\phi}$ derive from an Airy stress function $\phi$ eliminates the first equation and recasts the second into Pucher’s equation [1]
@@
Now if a displacement $(\mathbf{u}\equiv(u_{1},u_{2}),w)$ is isometric, then the infinitesimal membrane strain
@@
[…]
The analogy is well-known. Calladine in particular noted that $\mathbf{H}_{z}:\operatorname{cof}\mathbf{H}_{\psi}=0$ describes the absence of transverse loads when the test field $\psi$ is an Airy stress function, and the conservation of Gauss curvature when the test field $\psi$ is a transverse deflection [6]. Here, the novelty shall be in completing the analogy with integral formulas for average strains and stresses. In particular, for a periodic shell, let
@@
be the displacement field due to an effective membrane strain $\mathbf{E}$ and an effective bending strain $\bm{\chi}$. This is but the displacement that would reign in a homogeneous plate of uniform thickness, up to a correction allowing to relax some of the strain energy in function of the shell’s profile and composition. For a brief overview of this ansatz and of the elasticity of thin periodic plates, see [37]. The effective strains $(\mathbf{E},\bm{\chi})$ can be recovered from the local strain measures $(\bm{\varepsilon},\mathbf{H}_{w})$ by averaging. If the deformation is isometric, the averages $\left\langle{\cdot}\right\rangle$ can be written as
@@
thanks to ansatz (4), the divergence theorem and the vanishing of boundary terms by periodicity. Unless otherwise indicated, the subscript “oo” will be used to denote isometric strains and will be dropped when $\mathbf{E}$ and $\bm{\chi}$ are meant to be general strains. Similarly, for a statically admissible periodic membrane stress $\bm{\sigma}$ of Airy stress $\phi$, the effective membrane and bending stresses are given by the resultants
@@
Thus, by comparing (5) and (6), it comes that the static-geometric analogy further matches $\operatorname{cof}\bm{\chi}_{o}$ and $\bm{\Sigma}$ as well as $\operatorname{cof}\mathbf{E}_{o}$ and $\mathbf{M}$. Overall, the analogy provides an isomorphism
@@
between the spaces of effective isometric deformations and of effective admissible stresses, where the $\operatorname{cof}$ operator acts as a $90^{\circ}$ rotation; see Figure 1. Note that isometries always refer to there being no stretching of the shell’s midsurface, i.e., to the vanishing of local strains $\bm{\varepsilon}$. In particular, an effective deformation $(\mathbf{E},\bm{\chi})$ is isometric if it is the outcome of a deflection with vanishing local strains.
The counting rule
What is lacking here, to apply arguments à la Maxwell-Calladine [7], is to identify the spaces of effective admissible stresses and isometric strains with the image and kernel of some structural matrix. In the present context, the appropriate matrix turns out to be the matrix of effective elasticity moduli $\mathbf{A}$. To define it, consider the minimization of the average membrane strain energy density $\left\langle{\bm{\varepsilon}:\mathbf{c}:\bm{\varepsilon}}\right\rangle$ over the periodic corrections of ansatz (4) for prescribed $(\mathbf{E},\bm{\chi})$. Therein, $\mathbf{c}$ is the shell’s membrane stiffness tensor. This is a partial minimization of a positive semi-definite quadratic form and results in another positive semi-definite quadratic form
[…]
whose coefficients are identified as $\mathbf{A}$. The membrane strain energy conveniently ignores bending so that isometric deformations are zero modes of $\mathbf{A}$. In other words, for the purposes of counting, the shell is modeled as a membrane locally, and as a plate (Kirchhoff-Love) globally.
To conclude, the symmetry of $\mathbf{A}$ provides an orthogonal decomposition
$\text{Im}\mathbf{A}\oplus\text{Ker}\mathbf{A}=S^{2}\times S^{2}$ (9)
of the space of pairs of symmetric second order tensors $S^{2}\times S^{2}$. Then, the claim
$\text{Im}\mathbf{A}=\{(\bm{\Sigma},\mathbf{M})\},\quad\text{Ker}\mathbf{A}=\{(\mathbf{E}_{o},\bm{\chi}_{o})\}$ (10)
recasts the same decomposition into
$\{(\mathbf{\Sigma},\mathbf{M})\}\oplus\{(\mathbf{E}_{o},\bm{\chi}_{o})\}=S^{2}\times S^{2}.$ (11)
The analogy, on the other hand, showed that these spaces are isomorphic:
$\{(\mathbf{\Sigma},\mathbf{M})\}\cong\{(\mathbf{E}_{o},\bm{\chi}_{o})\}.$ (12)
Taking the dimensions then yields the counting rule
$\dim\{(\mathbf{E}_{o},\bm{\chi}_{o})\}=\frac{1}{2}\dim(S^{2}\times S^{2})=3.$ (13)
The Hill-Mandel lemma
However intuitive, claim (10) remains heuristic and warrants some attention. For instance, it is not quite obvious that admissible effective stresses $(\bm{\Sigma},\mathbf{M})$ can all be realized as the response of the shell to some imposed deflection for these responses could reasonably depend, say, on the composition of the shell. This turns out not to be the case thanks to a profound general result of homogenization theory known as the Hill-Mandel lemma [15]. The lemma states that the virtual stress-strain work is the same if computed using the local fields or the effective fields. Namely, for any admissible periodic $\bm{\sigma}$ and any admissible $(\mathbf{u},w)$ in the sense of (4), one has
@@
The proof is straightforward and employs the divergence theorem in conjunction with periodicity (4), admissibility (1) and the compatibility conditions (3); the Supplemental Material contains a proof. Most importantly, the lemma ensures that the static definition of stresses, by the resultants (6), is consistent with the thermodynamic definition
$\bm{\Sigma}=\frac{1}{2}\frac{\partial}{\partial\mathbf{E}}\left\langle{\bm{\varepsilon}:\mathbf{c}:\bm{\varepsilon}}\right\rangle,\quad\mathbf{M}=\frac{1}{2}\frac{\partial}{\partial\bm{\chi}}\left\langle{\bm{\varepsilon}:\mathbf{c}:\bm{\varepsilon}}\right\rangle.$ (15)
That is: the image of $\mathbf{A}$ is spanned by the resultants $(\bm{\Sigma},\mathbf{M})$ of membrane stresses $\bm{\sigma}\equiv\mathbf{c}:\bm{\varepsilon}$. A full rigorous proof of the identities in (10) is detailed in the supplemental material.
Microstructure-independent relationships for effective membrane moduli
The static-geometric isomorphism endows the space of effective isometric deformations with a natural symplectic structure: for any pair of modes $(\mathbf{E}_{i},\bm{\chi}_{i})_{i=1,2}$ in that space
@@
This is a restatement of the Hill-Mandel lemma (14) with the isomorphism (7) in mind and with $\bm{\varepsilon}=\mathbf{0}$ given that the involved strains are isometric. Such a relationship was recently derived using purely geometric tools in cases where either mode is a pure membrane mode, e.g., $\bm{\chi}_{1}=\mathbf{0}$, so that
$\mathbf{E}_{1}:\operatorname{cof}\bm{\chi}_{2}=0;$ (17)
see [31, 33]. In other words, the space of effective isometric deformations that are pure membrane modes is orthogonal to the space of isometric bending strains (rotated through $90^{\circ}$). This orthogonality, alone, shows that there can be no more than three effective isometric deformations. Here, with the help of the static-geometric analogy, and the minimum energy formulation of the homogenization problem, existence of exactly three modes is proven.
Equation (17) can be recast as an equality between in-plane and out-of-plane Poisson’s coefficients. This has been explored by many authors in many cases; see, e.g., [38, 42, 29, 36, 30, 43]. Thanks to the isomorphism constructed by the analogy, it is possible to draw a whole series of exact relationships among the effective moduli of a periodic membrane, all summarized into
@@
This relationship is more robust than the counting rule strictly speaking. Indeed, one can reasonably assume that the effective stiffness tensor $\mathbf{A}$ depends continuously on the local membrane stiffness $\mathbf{c}$ which can then be taken to zero at places so as to create holes. By contrast, the above relationship can certainly fail in the presence of handles, in which case $\mathbf{A}$ can become full rank (e.g., as in foldcore sandwich panels [22] and in Figure 2a).
@@
The role of topology
The topology of the shell plays a crucial role that remained implicit in the presentation above and that should be highlighted. On one hand, suppose the shell had handles, then there would be topological obstructions to integrating an admissible stress $\bm{\sigma}$ into an Airy stress function $\phi$. The static-geometric analogy, for averages, will then only encapsulate a subspace of admissible states of stress. The rule needs to be weakened into
$\dim\{(\mathbf{E}_{o},\bm{\chi}_{o})\}\leq 3\quad\text{(with handles, no holes)}.$ (19)
On the other hand, suppose the shell had holes, then there would be extra admissibility constraints for stresses in the form of zero tractions at boundaries. The analogy fails by leaving out isometries that do not reproduce this condition, i.e., isometries that cannot be extended over the holes or, equivalently, that deform the holes. The rule then weakens into
$\dim\{(\mathbf{E}_{o},\bm{\chi}_{o})\}\geq 3\quad\text{(with holes, no handles)}.$ (20)
Figure 2: Influence of topology beyond simple connectivity: (a) example of a surface with handles (three per unit cell) that has no effective isometries; (b) example of a surface with holes (one connected component per unit cell) that has six effective isometries. Both surfaces are a collection of cylinders with three sets of parallel axes (one of which is drawn) but with different topologies. Each family of closed cylinders in (a) can resist one uniaxial membrane stress and one uniaxial bending stress meaning, by the static-geometric analogy, that it admits no effective isometries. Each family of open cylinders in (b) can be isometrically flattened (following straight arrows) or bent (following circular arrows) to accommodate six (two modes, three axes) effective isometries.
To illustrate the way in which topology influences the count, consider the example of a “thickened” triangular truss as depicted on Figure 2a. This surface is not simply connected (with three handles per unit cell) and, quite intuitively, has no effective isometric deformations. At the other extreme, consider the “cork-inspired” honeycomb of Figure 2b. The wiggles, being cylindrical, allow for uniaxial stretching in three directions orthogonal to their axes as well as for uniaxial bending in three directions about their axes. Thus, this surface admits six effective isometries, i.e., the maximum possible. Clearly, this surface is not simply connected (with one connected boundary component per unit cell). Simply connected surfaces, as shown here, are precisely in the middle of this spectrum with three zero modes.
The unlikelihood of a conjecture
The three effective isometries of a simply connected surface generally have both membrane ($\mathbf{E}_{o}\neq\mathbf{0}$) and bending ($\bm{\chi}_{o}\neq\mathbf{0}$) components. If not, then we say that they are a pure membrane (if $\bm{\chi}_{o}=\mathbf{0}$) and a pure flexure (if $\mathbf{E}_{o}=\mathbf{0}$). The plane is an example of a surface that has no pure membrane modes and three pure flexure modes. The simple corrugation is an example with a single pure membrane mode and two pure flexure modes; see Figure 3a, b. Beyond that, there are no trivial examples with two or three pure membrane modes. It is not known if there is a fundamental limit on the number of pure membrane modes, but this appears to be unlikely. Here, the unit cells of surfaces with two and three approximately pure membrane modes are produced by a numerical minimization procedure. To do so, the surface is discretized into a truss and the $6\times 6$ effective elasticity matrix $\mathbf{A}$ is computed under periodic boundary conditions. The effective moduli corresponding to the target effective isometries are then weighed against the other moduli. The ratio, called $r$, provides an objective function that is minimized by a standard routine. It is given by
$r=\frac{\left\lVert[A_{ij}]_{i\in I,j\in\{1\dots 6\}}\right\rVert}{\left\lVert[A_{ij}]_{i,j=\{1\dots 6\}\setminus I}\right\rVert}.$ (21)
To obtain two pure membrane modes and a single pure flexure mode (a twist), we set $I=\{1,2,6\}$ corresponding to two uniaxial stretches and one twist. To obtain three pure membrane modes, we set $I=\{1,2,3\}$ corresponding to all membrane modes. Either way, the optimization parameter is the elevation map above a fixed triangular lattice, centered and normalized to avoid spurious elevations:
$\min r\text{ over }\{z\text{ periodic}\mid\left\langle{z}\right\rangle=0,\,\max z=z_{M}\}.$ (22)
See  [^1] Code is available at https://github.com/nassarh/discreteMembranes for implementation details. Figure 3 shows the results. The optimal geometries achieve a relative reduction in the moduli by a factor of $r\approx 10^{-8}$.
The above results strongly suggest that there are no restrictions on what effective isometries are realizable, beyond the implications of the simply connected topology, i.e., the count and the symplectic identity (16). In other words, it is suggested that any 3D subspace of $S^{2}\times S^{2}$, whose elements satisfy (16), is realizable as the space of effective isometries of some simply connected surface. In the jargon of symplectic linear algebra, such subspaces are called Lagrangians. Interestingly, Lagrangian subspaces are maximal: if the surface were to gain additional isometries, by weakening its topology, then the gained isometry will not satisfy (16). The set of all Lagrangian subspaces, i.e., of all candidate spaces of effective isometries, is the Lagrangian Grassmannian and is known to be a manifold of dimension six (in the present case). This means that the subspace of effective isometries depends continuously on a set of six coordinates. These coordinates are the $3\times 6$ components of a set of three zero modes $(\mathbf{E},\bm{\chi})$, minus the three constraints provided by identity (16), and taken modulo the 9-dimensional general linear group of $\mathbb{R}^{3}$ corresponding to transformations that shuffle and combine the zero modes but leave the subspace as is. See [25] for more details on Lagrangian Grassmannians.
Extensions
It is worthwhile to discuss how the results generalize to cases where the shell’s midsurface is piecewise smooth, i.e., with jumps in the tangent plane along crease lines (or crease curves). In these cases, admissible stresses and compatible strains can be discontinuous, albeit with precise jumps that guarantee local force balance and displacement continuity. Specifically, normal stresses and tangent strains remain continuous across crease lines. The normal and tangent directions being at $90^{\circ}$ to one another, the static-geometric analogy remains valid. The Hill-Mandel lemma remains valid as well, in the same fashion that it is valid in the presence of perfect material interfaces. See [33, 5, 2] for more detail. By the analogy and the lemma, the count holds. For non-periodic, statistically homogeneous, shells such as wrinkled or rough thin material surfaces, the Hill-Mandel lemma is known to remain valid in the limit of fast fluctuations, i.e., when the size of the shell is infinite relative to the representative volume element. Mathematical literature refers to this result as the $\mathrm{div}$-$\mathrm{curl}$ lemma; see [13]. Passing to the limit of fast fluctuations relieves the lemma, and therefore the result, of the burden of periodic boundary conditions.
Figure 3: Surfaces with (a) zero, (b) one, (c) two and (d) three pure membrane modes. Surfaces in (c) and (d) are obtained using the numerical optimization procedure described in the text; their count is not known to be exact albeit strongly suggested by the numerics.
Conclusion